OpenStudy (anonymous):

Hello, I'm trying to understand d/dx(2/sqrt(x)) please help.

OpenStudy (anonymous):

\[so, d/dx \left[ 2/\sqrt{x} \right]\]

OpenStudy (anonymous):

= 2 d/dx [x^(-1/2)] = 2(- 1/2)[ x^(-1/2 -1)] =- x^(-3/2)

OpenStudy (anonymous):

Thank you! I've posted one more question about the same subject.